The Greek Connection had sales of

$ 30.2

million and a cost of goods sold of

$ 12.1

a. Calculate The Greek​ Connection's net working capital in 2015.

b. Calculate the cash conversion cycle of The Greek Connection in 2015.

c. The industry average accounts receivable days is 30 days. What would the cash conversion cycle for The Greek Connection have been in 2015 had it matched the industry average for accounts receivable​ days?
Answer #1

Answer to Part a.

Net Working Capital = Current Assets – Current Liabilities
Net Working Capital = $6,832 - $3,879
Net Working Capital = $2,953

Answer to Part b.

Cash Conversion Cycle = Days Inventory Outstanding + Days Sales Outstanding - Days Payable Outstanding
Days Inventory Outstanding = Inventory / Cost of Goods Sold * 365
Days Inventory Outstanding = 1,274,000 / 12,100,000 * 365
Days Inventory Outstanding = 38.43 days

Days Sales Outstanding = Accounts Receivable / Net Credit Sales * 365
Days Sales Outstanding = 3,527,000 / 30,200,000 * 365
Days Sales Outstanding = 42.63 days

Days Payable Outstanding = Accounts Payable / Cost of Goods Sold * 365
Days Payable Outstanding = 1,659,000 / 12,100,000 * 365
Days Payable Outstanding = 50.04 days

Cash Conversion Cycle = 38.43 + 42.63 – 50.04
Cash Conversion Cycle = 31.02 days or 31 days

Answer to Part c.

Cash Conversion Cycle = Days Inventory Outstanding + Days Sales Outstanding - Days Payable Outstanding
Cash Conversion Cycle = 38.43 + 30.00 – 50.04
Cash Conversion Cycle = 18.39 or 18 days